OpenStudy (anonymous):

A park in a subdivision is triangular-shaped. Two adjacent sides of the park are 533 feet and 525 feet. The angle between the sides is 53°. Find the area of the park to the nearest square foot.

hero (hero):

Wow this problem looks very familiar.

Directrix (directrix):

d. 12,415 yd --> ANS Given any triangle, the area = 1/2 * product of any two sides * sine of included angle. Because the answer is required to be in square yards, divide the two lengths of 533 and 525 by 3 before multiplying. A = (1/2) (177.67) ( 175) (sin(53)) A = 12415.7 square yards (approx)

hero (hero):

Directrix, it did not need to be converted to yards